3.1 The average dividend yield of a random sample of 25 JSE-listed companies this year was found
to be 14.5%, with a sample standard deviation of 3.4%. Assume that dividend yields are
normally distributed.
3.1.1 Calculate, with 90% confidence, the actual mean dividend yield of all JSE-listed
companies this year. Interpret the finding. (6)
3.1.2 Calculate, with 95% confidence, the actual mean dividend yield of all JSE-listed
companies this year. Compare the interval with the one calculated in 3.1.1.
(12
1
Expert's answer
2021-09-17T03:44:15-0400

3.1.1

The critical value for "\alpha=0.1" is "z_c=z_{1-\alpha/2}=1.6449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(0.145-1.6449\times\dfrac{0.034}{\sqrt{25}}, 0.145+1.6449\times\dfrac{0.034}{\sqrt{25}})"

"=(0.1338, 0.1562)"

Therefore, based on the data provided, the 90 % confidence interval for the population mean is "0.1338<\mu<0.1562," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(0.1338, 0.1562),"

"(13.38\%, 15.62\%)"


3.1.2

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(0.145-1.95\times\dfrac{0.034}{\sqrt{25}}, 0.145+1.96\times\dfrac{0.034}{\sqrt{25}})"

"=(0.1317, 0.1583)"

Therefore, based on the data provided, the 95 % confidence interval for the population mean is "0.1317<\mu<0.1583," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(0.1317, 0.1583),"

"(13.17\%, 15.83\%)"


We find that the 95% confidence level has a wider interval as compared to the 90% confidence interval.

Increasing the confidence level increases the error bound, making the confidence interval wider. Decreasing the confidence level decreases the error bound, making the confidence interval narrower.